A marble is drawn at random from a bag containing 2 blue marbles, 3 yellow marbles, and 7 orange marbles. What is the probability that the marble drawn is NOT yellow?

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

there's no need to use CAPS on NOT.
--------
Do you think we can't READ?

Answer by greenestamps(13200)



2+3+7 = 12 marbles; 2+7 = 9 are not yellow.

Answer: P(not yellow) = 9/12 = 3/4